Linux Shadow Password HOWTO

Michael H. Jackson, mhjack@tscnet.com v1.3, 3 April 1996

This document aims to describe how to obtain, install, and configure the Linux password Shadow Suite. It also discusses obtaining, and reinstalling other software and network daemons that require access to user passwords. This other software is not actually part of the Shadow Suite, but these programs will need to be recompiled to support the Shadow Suite. This document also contains a programming example for adding shadow support to a program. Answers to some of the more frequently asked questions are included near the end of this document.

Contents

1 Introduction. 3 1.1 Changes from the previous release. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 New versions of this document. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Feedback. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Why shadow your passwd file? 4 2.1 Why you might NOT want to shadow your passwd file. . . . . . . . . . . . . . . . . . . . . . 6 2.2 Format of the /etc/passwd file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Format of the shadow file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4 Review of crypt(3). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 Getting the Shadow Suite. 9 3.1 History of the Shadow Suite for Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.2 Where to get the Shadow Suite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3 What is included with the Shadow Suite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Compiling the programs. 11 4.1 Unpacking the archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Configuring with the config.h file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3 Making backup copies of your original programs. . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.4 Running make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 Installing 13 5.1 Have a boot disk handy in case you break anything. . . . . . . . . . . . . . . . . . . . . . . . 13 5.2 Removing duplicate man pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.3 Running make install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.4 Running pwconv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.5 Renaming npasswd and nshadow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

6 Other programs you may need to upgrade or patch 15 6.1 Slackware adduser program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 6.2 The wu ftpd Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 6.3 Standard ftpd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.4 pop3d (Post Office Protocol 3) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.5 xlock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.6 xdm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.7 sudo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.8 imapd (E-Mail pinepackage) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.9 pppd (Point-to-Point Protocol Server) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

7 Putting the Shadow Suite to use. 21

7.1 Adding, Modifying, and deleting users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.1.1 useradd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.1.2 usermod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.1.3 userdel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

7.2 The passwd command and passwd aging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.3 The login.defs file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.4 Group passwords. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.5 Consistency checking programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

7.5.1 pwck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.5.2 grpck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

7.6 Dial-up passwords. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

8 Adding shadow support to a C program 30 8.1 Header files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.2 libshadow.a library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.3 Shadow Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.4 Shadow Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

8.5 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

9 Frequently Asked Questions. 36

10 Copyright Message. 36

11 Miscellaneous and Acknowledgments. 37

1 Introduction.

This is the Linux Shadow-Password-HOWTO. This document describes why and how to add shadow password support on a Linux system. Some examples of how to use some of the Shadow Suite’s features is also included.

When installing the Shadow Suite and when using many of the utility programs, you must be logged in as root. When installing the Shadow Suite you will be making changes to system software, and it is highly recommended that you make backup copies of programs as indicated. I also recommend that you read and understand all the instructions before you begin.

1.1 Changes from the previous release.

Additions:
        Added a sub-section on why you might not want to install shadow
        Added a sub-section on updating the xdm program
        Added a section on how to put Shadow Suite features to work
        Added a section containing frequently asked questions
Corrections/Updates:
        Corrected html references on Sunsite
        Corrected section on wu-ftp to reflect adding -lshadow to the Makefile
        Corrected minor spelling and verbiage errors
        Changed section on wu-ftpd to support ELF
        Updated to reflect security problems in various login programs
        Updated to recommend the Linux Shadow Suite by Marek Michalkiewicz

1.2 New versions of this document.

The latest released version of this document can always be retrieved by anonymous FTP from:

sunsite.unc.edu

/pub/Linux/docs/HOWTO/Shadow-Password-HOWTO

or:

/pub/Linux/docs/HOWTO/other-formats/Shadow-Password-HOWTO{-html.tar,ps,dvi}.gz

or  via  the  World  Wide  Web  from  the

Linux Documentation Project Web Server <http://sunsite.unc.edu/mdw/linux.html>, at page: Shadow-

Password-HOWTO <http://sunsite.unc.edu/linux/HOWTO/Shadow-Password-HOWTO.html> or directly

from me, <mhjack@tscnet.com>. It will also be posted to the newsgroup: comp.os.linux.answers

This document is now packaged with the Shadow-YYDDMM packages.

1.3 Feedback.

Please send any comments, updates, or suggestions to me: Michael H. Jackson <mhjack@tscnet.com¿ The sooner I get feedback, the sooner I can update and correct this document. If you find any problems with it, please mail me directly as I very rarely stay up-to-date on the newsgroups.

2 Why shadow your passwd file?

By default, most current Linux distributions do not contain the Shadow Suite installed. This includes Slackware 2.3, Slackware 3.0, and other popular distributions. One of the reasons for this is that the copyright notices in the original Shadow Suite were not clear on redistribution if a fee was charged. Linux uses a GNU Copyright (sometimes refereed to as a Copyleft) that allows people to package it into a convenient package (like a CD-ROM distribution) and charge a fee for it.

The current maintainer of the Shadow Suite, Marek Michalkiewicz <marekm@i17linuxb.ists.pwr.wroc.pl¿ received the source code from the original author under a BSD style copyright that allowed redistribution. Now that the copyright issues are resolved, it is expected that future distributions will contain password shadowing by default. Until then, you will need to install it yourself.

If you installed your distribution from a CD-ROM, you may find that, even though the distribution did not have the Shadow Suite installed, some of the files you need to install the Shadow Suite may be on the CD-ROM.

However, Shadow Suite versions 3.3.1, 3.3.1-2, and shadow-mk all have security problems with their login program and several other suid root programs that came with them, and should no longer be used.

All of the necessary files may be obtained via anonymous FTP or through the World Wide Web.

On a Linux system without the Shadow Suite installed, user information including passwords is stored in the /etc/passwd file. The password is stored in an encrypted format. If you ask a cryptography expert, however, he or she will tell you that the password is actually in an encoded rather than encrypted format because when using crypt(3), the text is set to null and the password is the key. Therefore, from here on, I will use the term encoded in this document.

The algorithm used to encode the password field is technically referred to as a one way hash function. This is an algorithm that is easy to compute in one direction, but very difficult to calculate in the reverse direction. More about the actual algorithm used can be found in section 2.4 or your crypt(3) manual page.

When a user picks or is assigned a password, it is encoded with a randomly generated value called the salt. This means that any particular password could be stored in 4096 different ways. The salt value is then stored with the encoded password.

When a user logs in and supplies a password, the salt is first retrieved from the stored encoded password. Then the supplied password is encoded with the salt value, and then compared with the encoded password. If there is a match, then the user is authenticated.

It is computationally difficult (but not impossible) to take a randomly encoded password and recover the original password. However, on any system with more than just a few users, at least some of the passwords will be common words (or simple variations of common words).

System crackers know all this, and will simply encrypt a dictionary of words and common passwords using all possible 4096 salt values. Then they will compare the encoded passwords in your /etc/passwd file with their database. Once they have found a match, they have the password for another account. This is referred to as a dictionary attack, and is one of the most common methods for gaining or expanding unauthorized access to a system.

If you think about it, an 8 character password encodes to 4096 * 13 character strings. So a dictionary of say 400,000 common words, names, passwords, and simple variations would easily fit on a 4GB hard drive. The attacker need only sort them, and then check for matches. Since a 4GB hard drive can be had for under $1000.00, this is well within the means of most system crackers.

Also, if a cracker obtains your /etc/passwd file first, they only need to encode the dictionary with the salt values actually contained in your /etc/passwd file. This method is usable by your average teenager with a couple of hundred spare Megabytes and a 486 class computer.

Even without lots of drive space, utilities like crack(1) can usually break at least a couple of passwords on a system with enough users (assuming the users of the system are allowed to pick their own passwords).

The /etc/passwd file also contains information like user ID’s and group ID’s that are used by many system programs. Therefore, the /etc/passwd file must remain world readable. If you were to change the /etc/passwd file so that nobody can read it, the first thing that you would notice is that the ls -l command now displays user ID’s instead of names!

The Shadow Suite solves the problem by relocating the passwords to another file (usually /etc/shadow). The /etc/shadow file is set so that it cannot be read by just anyone. Only root will be able to read and write to the /etc/shadow file. Some programs (like xlock) don’t need to be able to change passwords, they only need to be able to verify them. These programs can either be run suid root or you can set up a group shadow that is allowed read only access to the /etc/shadow file. Then the program can be run sgid shadow.

By moving the passwords to the /etc/shadow file, we are effectively keeping the attacker from having access to the encoded passwords with which to perform a dictionary attack.

Additionally, the Shadow Suite adds lots of other nice features:

• A configuration file to set login defaults (/etc/login.defs)

• Utilities for adding, modifying, and deleting user accounts and groups

• Password aging and expiration • Account expiration and locking

• Shadowed group passwords (optional)

• Double length passwords (16 character passwords) NOT RECOMMENDED

• Better control over user’s password selection

• Dial-up passwords

• Secondary authentication programs NOT RECOMMENDED Installing the Shadow Suite contributes toward a more secure system, but there are many other things that can also be done to improve the security of a Linux system, and there will eventually be a series of Linux Security HOWTO’s that will discuss other security measures and related issues. For current information on other Linux security issues, including warnings on known vulnerabilities see the

Linux Security home page. <http://bach.cis.temple.edu/linux/linux-security/>

2.1 Why you might NOT want to shadow your passwd file.

There are a few circumstances and configurations in which installing the Shadow Suite would NOT be a good idea:

• The machine does not contain user accounts.

• Your machine is running on a LAN and is using NIS (Network Information Services) to get or supply user names and passwords to other machines on the network. (This can actually be done, but is beyond the scope of this document, and really won’t increase security much anyway)

• Your machine is being used by terminal servers to verify users via NFS (Network File System), NIS, or some other method.

• Your machine runs other software that validates users, and there is no shadow version available, and you don’t have the source code.

2.2 Format of the /etc/passwd file

A non-shadowed /etc/passwd file has the following format:

     username:passwd:UID:GID:full_name:directory:shell

Where:

username

The user (login) name

passwd

The encoded password

UID

Numerical user ID

GID

Numerical default group ID

full name

The user’s full name - Actually this field is called the GECOS (General Electric Comprehensive Operating System) field and can store information other than just the full name. The Shadow commands and manual pages refer to this field as the comment field.

directory

User’s home directory (Full pathname)

shell

User’s login shell (Full Pathname)

For example:

     username:Npge08pfz4wuk:503:100:Full Name:/home/username:/bin/sh

Where Np is the salt and ge08pfz4wuk is the encoded password. The encoded salt/password could just as easily have been kbeMVnZM0oL7I and the two are exactly the same password. There are 4096 possible encodings for the same password. (The example password in this case is ’password’, a really bad password). Once the shadow suite is installed, the /etc/passwd file would instead contain:

     username:x:503:100:Full Name:/home/username:/bin/sh

The x in the second field in this case is now just a place holder. The format of the /etc/passwd file really didn’t change, it just no longer contains the encoded password. This means that any program that reads the /etc/passwd file but does not actually need to verify passwords will still operate correctly. The passwords are now relocated to the shadow file (usually /etc/shadow file).

2.3 Format of the shadow file

The /etc/shadow file contains the following information:

     username:passwd:last:may:must:warn:expire:disable:reserved

Where:

username

The User Name

passwd

The Encoded password

last

Days since Jan 1, 1970 that password was last changed

may

Days before password may be changed

must

Days after which password must be changed

warn

Days before password is to expire that user is warned

expire

Days after password expires that account is disabled

disable

Days since Jan 1, 1970 that account is disabled

reserved

A reserved field

The previous example might then be:

     username:Npge08pfz4wuk:9479:0:10000::::

2.4 Review of crypt(3).

From the crypt(3) manual page: "crypt is the password encryption function. It is based on the Data Encryption Standard algorithm with variations intended (among other things) to discourage use of hardware implementations of a key search. T he key is a user’s typed password. T heencodedstringisallNULLs T he salt is a two-character string chosen from the set a − zA − Z0 − 9./. This string is used to perturb the algorithm in one of 4096 different ways. By taking the lowest 7 bits of each character of the key, a 56-bit key is obtained. This 56-bit key is used to encrypt repeatedly a constant string (usually a string consisting of all zeros). The returned value points to the encrypted password, a series of 13 printable ASCII characters (the first two characters represent the salt itself). The return value points to static data whose content is overwritten by each call.

Warning: The key space consists of 2**56 equal 7.2e16 possible values. Exhaustive searches of this key space are possible using massively parallel computers. Software, such as crack(1), is available which will search the portion of this key space that is generally used by humans for passwords. Hence, password selection should, at minimum, avoid common words and names. The use of a passwd(1) program that checks for crackable passwords during the selection process is recommended.

The DES algorithm itself has a few quirks which make the use of the crypt(3) interface a very poor choice for anything other than password authentication. If you are planning on using the crypt(3) interface for a cryptography project, don’t do it: get a good book on encryption and one of the widely available DES libraries."

Most Shadow Suites contain code for doubling the length of the password to 16 characters. Experts in des recommend against this, as the encoding is simply applied first to the left half and then to the right half of the longer password. Because of the way crypt works, this may make for a less secure encoded password then if double length passwords were not used in the first place. Additionally, it is less likely that a user will be able to remember a 16 character password.

There is development work under way that would allow the authentication algorithm to be replaced with something more secure and with support for longer passwords (specifically the MD5 algorithm) and retain compatibility with the crypt method.

If you are looking for a good book on encryption, I recommend:

        "Applied Cryptography: Protocols, Algorithms, and Source Code in C"
        by Bruce Schneier <schneier@chinet.com>
        ISBN: 0-471-59756-2

3 Getting the Shadow Suite.

3.1 History of the Shadow Suite for Linux

DO NOT USE THE PACKAGES IN THIS SECTION, THEY HAVE SECURITY PROBLEMS

The original Shadow Suite was written by John F. Haugh II.

There are several versions that have been used on Linux systems:

• shadow-3.3.1 is the original.

• shadow-3.3.1-2 is Linux specific patch made by Florian La Roche <flla@stud.uni-sb.de¿ and contains some further enhancements.

• shadow-mk was specifically packaged for Linux.

The shadow-mk package contains the shadow-3.3.1 package distributed by John F. Haugh II with the shadow-3.3.1-2 patch installed, a few fixes made by Mohan Kokal <magnus@texas.net¿ that make installation a lot easier, a patch by Joseph R.M. Zbiciak for login1.c (login.secure) that eliminates the -f, -h security holes in /bin/login, and some other miscellaneous patches.

The shadow.mk package was the previously recommended package, but should be replaced due to a security problem with the login program.

There are security problems with Shadow versions 3.3.1, 3.3.1-2, and shadow-mk involving the login program. This login bug involves not checking the length of a login name. This causes the buffer to overflow causing crashes or worse. It has been rumored that this buffer overflow can allow someone with an account on the system to use this bug and the shared libraries to gain root access. I won’t discuss exactly how this is possible because there are a lot of Linux systems that are affected, but systems with these Shadow Suites installed, and most pre-ELF distributions without the Shadow Suite are vulnerable!

For more information on this and other Linux security issues, see the Linux Security home page (Shared

Libraries  and  login  Program  Vulnerability)

<http://bach.cis.temple.edu/linux/linux-security/Linux-Security-FAQ/Linux-telnetd.html>

3.2 Where to get the Shadow Suite.

The only recommended Shadow Suite is still in BETA testing, however the latest versions are safe in a production environment and don’t contain a vulnerable login program.

The package uses the following naming convention:

     shadow-YYMMDD.tar.gz

where YYMMDD is the issue date of the Suite.

This version will eventually be Version 3.3.3 when it is released from Beta testing, and is maintained by Marek Michalkiewicz <marekm@i17linuxb.ists.pwr.wroc.pl¿. It’s available as: shadow-current.tar.gz

<ftp://i17linuxb.ists.pwr.wroc.pl/pub/linux/shadow/shadow-current.tar.gz>.

The following mirror sites have also been established:

• ftp://ftp.icm.edu.pl/pub/Linux/shadow/shadow-current.tar.gz

• ftp://iguana.hut.fi/pub/linux/shadow/shadow-current.tar.gz

• ftp://ftp.cin.net/usr/ggallag/shadow/shadow-current.tar.gz

• ftp://ftp.netural.com/pub/linux/shadow/shadow-current.tar.gz

You should use the currently available version.

You should NOT use a version older than shadow-960129 as they also have the login security problem discussed above.

When this document refers to the Shadow Suite I am referring to the this package. It is assumed that this is the package that you are using.

For reference, I used shadow-960129 to make these installation instructions.

If you were previously using shadow-mk, you should upgrade to this version and rebuild everything that you originally compiled.

3.3 What is included with the Shadow Suite.

The Shadow Suite contains replacement programs for:

su, login, passwd, newgrp, chfn, chsh, and id

The package also contains the new programs:

chage, newusers, dpasswd, gpasswd, useradd, userdel, usermod, groupadd, groupdel,
groupmod, groups, pwck, grpck, lastlog, pwconv, and pwunconv

Additionally, the library: libshadow.a is included for writing and/or compiling programs that need to access user passwords. Also, manual pages for the programs are also included. There is also a configuration file for the login program which will be installed as /etc/login.defs.

4 Compiling the programs.

4.1 Unpacking the archive.

The first step after retrieving the package is unpacking it. The package is in the tar (tape archive) format and compressed using gzip, so first move it to /usr/src, then type:

     tar -xzvf shadow-current.tar.gz

This will unpack it into the directory: /usr/src/shadow-YYMMDD

4.2 Configuring with the config.h file

The first thing that you need to do is to copy over the Makefile and the config.h file:

     cd /usr/src/shadow-YYMMDD
     cp Makefile.linux Makefile
     cp config.h.linux config.h

You should then take a look at the config.h file. This file contains definitions for some of the configuration options. If you are using the recommended package, I recommend that you disable group shadow support for your first time around. By default shadowed group passwords are enabled. To disable these edit the config.h file, and change the #define SHADOWGRP to #undef SHADOWGRP. I recommend that you disable them to start with, and then if you really want group passwords and group administrators that you enable it later and recompile. If you leave it enabled, you must create the file /etc/gshadow. Enabling the long passwords option is NOT recommended as discussed above. Do NOT change the setting: #undef AUTOSHADOW The AUTOSHADOW option was originally designed so that programs that were shadow ignorant would still function. This sounds good in theory, but does not work correctly. If you enable this option, and the program runs as root, it may call getpwnam() as root, and later write the modified entry back to the /etc/passwd file (with the no-longer-shadowed password). Such programs include chfn and chsh. (You can’t get around this by swapping real and effective uid before calling getpwnam() because root may use chfn and chsh too.)

The same warning is also valid if you are building libc, it has a SHADOW COMPAT option which does the same thing. It should NOT be used! If you start getting encoded passwords back in your /etc/passwd file, this is the problem.

If you are using a libc version prior to 4.6.27, you will need to make a couple more changes to config.h and the Makefile. To config.h edit and change:

     #define HAVE_BASENAME

to:

     #undef HAVE_BASENAME

And then in the Makefile, change:

     SOBJS = smain.o env.o entry.o susetup.o shell.o \
             sub.o mail.o motd.o sulog.o age.o tz.o hushed.o
     SSRCS = smain.c env.c entry.c setup.c shell.c \
             pwent.c sub.c mail.c motd.c sulog.c shadow.c age.c pwpack.c rad64.c \
             tz.c hushed.c
     SOBJS = smain.o env.o entry.o susetup.o shell.o \
             sub.o mail.o motd.o sulog.o age.o tz.o hushed.o basename.o
     SSRCS = smain.c env.c entry.c setup.c shell.c \
             pwent.c sub.c mail.c motd.c sulog.c shadow.c age.c pwpack.c rad64.c \
             tz.c hushed.c basename.c

These changes add the code contained in basename.c which is contained in libc 4.6.27 and later.

4.3 Making backup copies of your original programs.

It would also be a good idea to track down and make backup copies of the programs that the shadow suite will replace. On a Slackware 3.0 system these are:

• /bin/su

• /bin/login

• /usr/bin/passwd • /usr/bin/newgrp

• /usr/bin/chfn

• /usr/bin/chsh

• /usr/bin/id

The BETA package has a save target in the Makefile, but it’s commented out because different distributions place the programs in different places.

You should also make a backup copy of your /etc/passwd file, but be careful to name it something else if you place it in the same directory so you don’t overwrite the passwd command.

4.4 Running make

You need to be logged as root to do most of the installation.

Run make to compile the executables in the package:

     make all

You may see the warning: rcsid defined but not used. This is fine, it just happens because the author is using a version control package.

5 Installing

5.1 Have a boot disk handy in case you break anything.

If something goes terribly wrong, it would be handy to have a boot disk. If you have a boot/root combination from your installation, that will work, otherwise see the Bootdisk-HOWTO

<http://sunsite.unc.edu/mdw/HOWTO/Bootdisk-HOWTO.html>, which describes how to make a bootable

disk.

5.2 Removing duplicate man pages

You should also move the manual pages that are about to be replaced. Even if you are brave enough install the Shadow Suite without making backups, you will still want to remove the old manual pages. The new manual pages won’t normally overwrite the old ones because the old ones are probably compressed.

You can use a combination of: man -aW command and locate command to locate the manual pages that need to be (re)moved. It’s generally easier to figure out which are the older pages before you run make install.

If you are using the Slackware 3.0 distribution, then the manual pages you want to remove are:

• /usr/man/man1/chfn.1.gz • /usr/man/man1/chsh.1.gz

• /usr/man/man1/id.1.gz

• /usr/man/man1/login.1.gz

• /usr/man/man1/passwd.1.gz

• /usr/man/man1/su.1.gz

• /usr/man/man5/passwd.5.gz

There may also be man pages of the same name in the /var/man/cat[1-9] subdirectories that should also be deleted.

5.3 Running make install

You are now ready to type: (do this as root)

     make install

This will install the new and replacement programs and fix-up the file permissions. It will also install the man pages.

This also takes care of installing the Shadow Suite include files in the correct places in /usr/include/shadow.

Using the BETA package you must manually copy the file login.defs to the /etc subdirectory and make sure that only root can make changes to it.

     cp login.defs /etc
     chmod 700 /etc/login.defs

This file is the configuration file for the login program. You should review and make changes to this file for your particular system. This is where you decide which tty’s root can login from, and set other security policy settings (like password expiration defaults).

5.4 Running pwconv

The next step is to run pwconv. This must also be done as root, and is best done from the /etc subdirectory:

     cd /etc
     /usr/sbin/pwconv

pwconv takes your /etc/passwd file and strips out the fields to create two files: /etc/npasswd and

/etc/nshadow.

A pwunconv program is also provided if you need to make a normal /etc/passwd file out of an /etc/passwd and /etc/shadow combination.

5.5 Renaming npasswd and nshadow

Now that you have run pwconv you have created the files /etc/npasswd and /etc/nshadow. These need to be copied over to /etc/passwd and /etc/shadow. We also want to make a backup copy of the original /etc/passwd file, and make sure only root can read it. We’ll put the backup in root’s home directory:

     cd /etc
     cp passwd ~passwd
     chmod 600 ~passwd
     mv npasswd passwd
     mv nshadow shadow

You should also ensure that the file ownerships and permissions are correct. If you are going to be using X-Windows, the xlock and xdm programs need to be able to read the shadow file (but not write it).

There are two ways that this can be done. You can set xlock to suid root (xdm is usually run as root anyway). Or you can make the shadow file owned by root with a group of shadow, but before you do this, make sure that you have a shadow group (look in /etc/group). None of the users on the system should actually be in the shadow group.

     chown root.root passwd
     chown root.shadow shadow
     chmod 0644 passwd
     chmod 0640 shadow

Your system now has the password file shadowed. You should now pop over to another virtual terminal and verify that you can login.

Really, do this now!

If you can’t, then something is wrong! To get back to a non-shadowed state, do the following the following:

     cd /etc
     cp ~passwd passwd
     chmod 644 passwd

You would then restore the files that you saved earlier to their proper locations.

6 Other programs you may need to upgrade or patch

Even though the shadow suite contains replacement programs for most programs that need to access passwords, there are a few additional programs on most systems that require access to passwords.

If you are running a Debian Distribution (or even if you are not), you can obtain Debian sources for the programs that need to be rebuild from: ftp://ftp.debian.org/debian/stable/source/

The remainder of this section discusses how to upgrade adduser, wu ftpd, ftpd, pop3d, xlock, xdm and sudo so that they support the shadow suite.

See the section 8 for a discussion on how to put shadow support into any other program that needs it (although the program must then be run SUID root or SGID shadow to be able to actually access the shadow file).

6.1 Slackware adduser program

Slackware distributions (and possibly some others) contain a interactive program for adding users called /sbin/adduser. A shadow version of this program can be obtained from ftp://sunsite.unc.edu/pub/Linux/ system/Admin/accounts/adduser.shadow-1.4.tar.gz.

I would encourage you to use the programs that are supplied with the Shadow Suite (useradd, usermod, and userdel) instead of the slackware adduser program. They take a little time to learn how to use, but it’s well worth the effort because you have much more control and they perform proper file locking on the

/etc/passwd and /etc/shadow file (adduser doesn’t).

See the section on 7 for more information.

But if you gotta have it, here is what you do:

     tar -xzvf adduser.shadow-1.4.tar.gz
     cd adduser
     make clean
     make adduser
     chmod 700 adduser
     cp adduser /sbin

6.2 The wu ftpd Server

Most Linux systems some with the wu ftpd server. If your distribution does not come with shadow installed, then your wu ftpd will not be compiled for shadow. wu ftpd is launched from inetd/tcpd as a root process. If you are running an old wu ftpd daemon, you will want to upgrade it anyway because older ones had a bug that would allow the root account to be compromised (For more info see the Linux security home page

<http://bach.cis.temple.edu/linux/linux-security/Linux-Security-FAQ/Linux-wu.ftpd-2.4-Update.html>).

Fortunately, you only need to get the source code and recompile it with shadow enabled.

If you are not running an ELF system, The wu ftp server can be found on Sunsite as wu-ftp-2.4-fixed.tar.gz

<ftp://sunsite.unc.edu/pub/Linux/system/Network/file-transfer/wu-ftpd-2.4-fixed.tar.gz>

Once you retrieve the server, put it in /usr/src, then type:

     cd /usr/src
     tar -xzvf wu-ftpd-2.4-fixed.tar.gz
     cd wu-ftpd-2.4-fixed
     cp ./src/config/config.lnx.shadow ./src/config/config.lnx
Then edit ./src/makefiles/Makefile.lnx, and change the line:
     LIBES    = -lbsd -support

to:

     LIBES    = -lbsd -support -lshadow

Now you are ready to run the build script and install:

     cd /usr/src/wu-ftpd-2.4-fixed
     /usr/src/wu-ftp-2.4.fixed/build lnx
     cp /usr/sbin/wu.ftpd /usr/sbin/wu.ftpd.old
     cp ./bin/ftpd /usr/sbin/wu.ftpd

This uses the Linux shadow configuration file, compiles and installs the server. On my Slackware 2.3 system I also had to do the following before running build:

     cd /usr/include/netinet
     ln -s in_systm.h in_system.h
     cd -

Problems  have  been  reported  compiling  this  package  under  ELF  sys-

tems, but the Beta version of the next release works fine. It can be found as wu-ftp-2.4.2-beta-10.tar.gz

<ftp://tscnet.com/pub/linux/network/ftp/wu-ftpd-2.4.2-beta-10.tar.gz>

Once you retrieve the server, put it in /usr/src, then type:

     cd /usr/src
     tar -xzvf wu-ftpd-2.4.2-beta-9.tar.gz
     cd wu-ftpd-beta-9
     cd ./src/config

Then edit config.lnx, and change:

     #undef SHADOW.PASSWORD

to:

     #define SHADOW.PASSWORD

Then,

     cd ../Makefiles

and edit the file Makefile.lnx and change:

     LIBES = -lsupport -lbsd # -lshadow

to:

     LIBES = -lsupport -lbsd -lshadow

Then build and install:

     cd ..
     build lnx
     cp /usr/sbin/wu.ftpd /usr/sbin/wu.ftpd.old
     cp ./bin/ftpd /usr/sbin/wu.ftpd

Note that you should check your /etc/inetd.conf file to make sure that this is where your wu.ftpd server really lives. It has been reported that some distributions place the server daemons in different places, and then wu.ftpd in particular may be named something else.

6.3 Standard ftpd

If you are running the standard ftpd server, I would recommend that you upgrade to the wu ftpd server. Aside from the known bug discussed above, it’s generally thought to be more secure.

If you insist on the standard one, or you need NIS support, Sunsite has ftpd-shadow-nis.tgz

<ftp://sunsite.unc.edu/pub/Linux/system/Network/file-transfer/ftpd-shadow-nis.tgz>

6.4 pop3d (Post Office Protocol 3)

If you need to support the third Post Office Protocol (POP3), you will need to recompile a pop3d program.

pop3d is normally run by inetd/tcpd as root.

There are two versions available from Sunsite: pop3d-1.00.4.linux.shadow.tar.gz

<ftp://sunsite.unc.edu/pub/Linux/system/Mail/pop/pop3d-1.00.4.linux.shadow.tar.gz>  and

pop3d+shadow+elf.tar.gz

<ftp://sunsite.unc.edu/pub/Linux/system/Mail/pop/pop3d+shadow+elf.tar.gz>

Both of these are fairly straight forward to install.

6.5 xlock

If you install the shadow suite, and then run X Windows System and lock the screen without upgrading your xlock, you will have to use CNTL-ALT-Fx to switch to another tty, login, and kill the xlock process (or use CNTL-ALT-BS to kill the X server). Fortunately it’s fairly easy to upgrade your xlock program.

If you are running XFree86 Versions 3.x.x, you are probably using xlockmore (which is a great screen-saver in addition to a lock). This package supports shadow with a recompile. If you have an older xlock, I recommend that you upgrade to this one.

xlockmore-3.5.tgz  is  available  at:
<ftp://sunsite.unc.edu/pub/Linux/X11/xutils/screensavers/xlockmore-3.7.tgz>

Basically, this is what you need to do:

Get the xlockmore-3.7.tgz file and put it in /usr/src unpack it:

     tar -xzvf xlockmore-3.7.tgz
Edit the file: /usr/X11R6/lib/X11/config/linux.cf, and change the line:
     #define HasShadowPasswd    NO
     to
     #define HasShadowPasswd    YES

Then build the executables:

     cd /usr/src/xlockmore
     xmkmf
     make depend
     make

Then move everything into place and update file ownerships and permissions:

     cp xlock /usr/X11R6/bin/
     cp XLock /var/X11R6/lib/app-defaults/
     chown root.shadow /usr/X11R6/bin/xlock
     chmod 2755 /usr/X11R6/bin/xlock
     chown root.shadow /etc/shadow
     chmod 640 /etc/shadow

Your xlock will now work correctly.

6.6 xdm

xdm is a program that presents a login screen for X-Windows. Some systems start xdm when the system is told to goto a specified run level (see /etc/inittab.

With the Shadow Suite install, xdm will need to be updated. Fortunately it’s fairly easy to upgrade your xdm program.

xdm.tar.gz is available at: <ftp://sunsite.unc.edu/pub/Linux/X11/xutils/xdm.tar.gz>

Get the xdm.tar.gz file and put it in /usr/src, then to unpack it:

     tar -xzvf xdm.tar.gz
Edit the file: /usr/X11R6/lib/X11/config/linux.cf, and change the line:
     #define HasShadowPasswd    NO
     to
     #define HasShadowPasswd    YES

Then build the executables:

     cd /usr/src/xdm
     xmkmf
     make depend
     make

Then move everything into place:

     cp xdm /usr/X11R6/bin/

xdm is run as root so you don’t need to change it file permissions.

6.7 sudo

The program sudo allows a system administrator to let users run programs that would normally require root access. This is handy because it lets the administrator limit access to the root account itself while still allowing users to do things like mounting drives.

sudo needs to read passwords because it verifies the users password when it’s invoked. sudo already runs SUID root, so accessing the /etc/shadow file is not a problem.

sudo  for  the  shadow  suite,  is  available  as  at:

<ftp://sunsite.unc.edu/pub/Linux/system/Admin/sudo-1.2-shadow.tgz>

Warning: When you install sudo your /etc/sudoers file will be replaced with a default one, so you need to make a backup of it if you have added anything to the default one. (you could also edit the Makefile and remove the line that copies the default file to /etc).

The package is already setup for shadow, so all that’s required is to recompile the package (put it in

/usr/src):
     cd /usr/src
     tar -xzvf sudo-1.2-shadow.tgz
     cd sudo-1.2-shadow
     make all
     make install

6.8 imapd (E-Mail pinepackage)

imapd is an e-mail server similar to pop3d. imapd comes with the Pine E-mail package. The documentation that comes with the package states that the default for Linux systems is to include support for shadow. However, I have found that this is not true. Furthermore, the build script / Makefile combination on this package is makes it very difficult to add the libshadow.a library at compile time, so I was unable to add shadow support for imapd.

If anyone has this figured out, please E-mail me, and I’ll include the solution here.

6.9 pppd (Point-to-Point Protocol Server)

The pppd server can be setup to use several types of authentication: Password Authentication Protocol (PAP) and Cryptographic Handshake Authentication Protocol (CHAP). The pppd server usually reads the password strings that it uses from /etc/ppp/chap-secrets and/or /etc/ppp/pap-secrets. If you are using this default behavior of pppd, it is not necessary to reinstall pppd.

pppd also allows you to use the login parameter (either on the command line, or in the configuration or options file). If the login option is given, then pppd will use the /etc/passwd file for the username and passwords for the PAP. This, of course, will no longer work now that our password file is shadowed. For pppd-1.2.1d this requires adding code for shadow support.

The example given in the next section is adding shadow support to pppd-1.2.1d (an older version of pppd).

pppd-2.2.0 already contains shadow support.

7 Putting the Shadow Suite to use.

This section discusses some of the things that you will want to know now that you have the Shadow Suite installed on your system. More information is contained in the manual pages for each command.

7.1 Adding, Modifying, and deleting users

The Shadow Suite added the following command line oriented commands for adding, modifying, and deleting users. You may also have installed the adduser program.

7.1.1 useradd

The useradd command can be used to add users to the system. You also invoke this command to change the default settings.

The first thing that you should do is to examine the default settings and make changes specific to your system:

     useradd -D
GROUP=1
HOME=/home
INACTIVE=0
EXPIRE=0
SHELL=
SKEL=/etc/skel

The defaults are probably not what you want, so if you started adding users now you would have to specify all the information for each user. However, we can and should change the default values.

On my system:

• I want the default group to be 100

• I want passwords to expire every 60 days

• I don’t want to lock an account because the password is expired

• I want to default shell to be /bin/bash

To make these changes I would use:

     useradd -D -g100 -e60 -f0 -s/bin/bash

Now running useradd -D will give:

GROUP=100
HOME=/home
INACTIVE=0
EXPIRE=60
SHELL=/bin/bash
SKEL=/etc/skel

Just in case you wanted to know, these defaults are stored in the file /etc/default/useradd.

Now you can use useradd to add users to the system. For example, to add the user fred, using the defaults, you would use the following:

     useradd -m -c "Fred Flintstone" fred

This will create the following entry in the /etc/passwd file:

     fred:*:505:100:Fred Flintstone:/home/fred:/bin/bash

And the following entry in the /etc/shadow file:

     fred:!:0:0:60:0:0:0:0

fred’s home directory will be created and the contents of /etc/skel will be copied there because of the -m switch.

Also, since we did not specify a UID, the next available one was used.

fred’s account is created, but fred still won’t be able to login until we unlock the account. We do this by changing the password.

     passwd fred

Changing password for fred
Enter the new password (minimum of 5 characters)
Please use a combination of upper and lower case letters and numbers.
New Password: *******
Re-enter new password: *******

Now the /etc/shadow will contain:

     fred:J0C.WDR1amIt6:9559:0:60:0:0:0:0

And fred will now be able to login and use the system. The nice thing about useradd and the other programs that come with the Shadow Suite is that they make changes to the /etc/passwd and /etc/shadow files atomically. So if you are adding a user, and another user is changing their password at the same time, both operations will be performed correctly.

You should use the supplied commands rather than directly editing /etc/passwd and /etc/shadow. If you were editing the /etc/shadow file, and a user were to change his password while you are editing, and then you were to save the file you were editing, the user’s password change would be lost.

Here is a small interactive script that adds users using useradd and passwd:

#!/bin/bash
#
# /sbin/newuser - A script to add users to the system using the Shadow
#                 Suite's useradd and passwd commands.
#
# Written my Mike Jackson <mhjack@tscnet.com> as an example for the Linux
# Shadow Password Howto.  Permission to use and modify is expressly granted.
#
# This could be modified to show the defaults and allow modification similar
# to the Slackware Adduser program.  It could also be modified to disallow
# stupid entries.  (i.e. better error checking).
#
##
#  Defaults for the useradd command
##
GROUP=100        # Default Group
HOME=/home       # Home directory location (/home/username)
SKEL=/etc/skel   # Skeleton Directory
INACTIVE=0       # Days after password expires to disable account (0=never)
EXPIRE=60        # Days that a passwords lasts
SHELL=/bin/bash  # Default Shell (full path)
##
#  Defaults for the passwd command
##

PASSMIN=0        # Days between password changes
PASSWARN=14      # Days before password expires that a warning is given
##
#  Ensure that root is running the script.
##
WHOAMI=`/usr/bin/whoami`
if [ $WHOAMI != "root" ]; then
        echo "You must be root to add news users!"
        exit 1
fi
##
#  Ask for username and fullname.
##
echo ""
echo -n "Username: "
read USERNAME
echo -n "Full name: "
read FULLNAME
#
echo "Adding user: $USERNAME."
#
# Note that the "" around $FULLNAME is required because this field is
# almost always going to contain at least on space, and without the "'s
# the useradd command would think that you we moving on to the next
# parameter when it reached the SPACE character.
#
/usr/sbin/useradd -c"$FULLNAME" -d$HOME/$USERNAME -e$EXPIRE \
        -f$INACTIVE -g$GROUP -m -k$SKEL -s$SHELL $USERNAME
##
#  Set password defaults
##
/bin/passwd -n $PASSMIN -w $PASSWARN $USERNAME >/dev/null 2>&1
##
#  Let the passwd command actually ask for password (twice)
##
/bin/passwd $USERNAME
##
#  Show what was done.
##
echo ""
echo "Entry from /etc/passwd:"
echo -n "   "
grep "$USERNAME:" /etc/passwd
echo "Entry from /etc/shadow:"

echo -n "   "
grep "$USERNAME:" /etc/shadow
echo "Summary output of the passwd command:"
echo -n "   "
passwd -S $USERNAME
echo ""

Using a script to add new users is really much more preferable than editing the /etc/passwd or /etc/shadow files directly or using a program like the Slackware adduser program. Feel free to use and modify this script for your particular system.

For more information on the useradd see the online manual page.

7.1.2 usermod

The usermod program is used to modify the information on a user. The switches are similar to the useradd program.

Let’s say that you want to change fred’s shell, you would do the following:

     usermod -s /bin/tcsh fred

Now fred’s /etc/passwd file entry would be change to this:

     fred:*:505:100:Fred Flintstone:/home/fred:/bin/tcsh

Let’s make fred’s account expire on 09/15/97:

     usermod -e 09/15/97 fred

Now fred’s entry in /etc/shadow becomes:

     fred:J0C.WDR1amIt6:9559:0:60:0:0:10119:0

For more information on the usermod command see the online manual page.

7.1.3 userdel

userdel does just what you would expect, it deletes the user’s account. You simply use:

     userdel -r username

The -r causes all files in the user’s home directory to be removed along with the home directory itself. Files located in other file system will have to be searched for and deleted manually.

If you want to simply lock the account rather than delete it, use the passwd command instead.

7.2 The passwd command and passwd aging.

The passwd command has the obvious use of changing passwords. Additionally, it is used by the root user to:

• Lock and unlock accounts (-l and -u)

• Set the maximum number of days that a password remains valid (-x)

• Set the minimum days between password changes (-n)

• Sets the number of days of warning that a password is about to expire (-w)

• Sets the number of days after the password expires before the account is locked (-i)

• Allow viewing of account information in a clearer format (-S)

For example, let look again at fred

     passwd -S fred
     fred P 03/04/96 0 60 0 0

This means that fred’s password is valid, it was last changed on 03/04/96, it can be changed at any time, it expires after 60 days, fred will not be warned, and and the account won’t be disabled when the password expires.

This simply means that if fred logs in after the password expires, he will be prompted for a new password at login.

If we decide that we want to warn fred 14 days before his password expires and make his account inactive 14 days after he lets it expire, we would need to do the following:

     passwd -w14 -i14 fred

Now fred is changed to:

     fred P 03/04/96 0 60 14 14

For more information on the passwd command see the online manual page.

7.3 The login.defs file.

The file /etc/login is the configuration file for the login program and also for the Shadow Suite as a whole.

/etc/login contains settings from what the prompts will look like to what the default expiration will be when a user changes his password.

The /etc/login.defs file is quite well documented just by the comments that are contained within it. However, there are a few things to note:

• It contains flags that can be turned on or off that determine the amount of logging that takes place.

• It contains pointers to other configuration files.

• It contains defaults assignments for things like password aging.

From the above list you can see that this is a rather important file, and you should make sure that it is present, and that the settings are what you desire for your system.

7.4 Group passwords.

The /etc/groups file may contain passwords that permit a user to become a member of a particular group. This function is enabled if you define the constant SHADOWGRP in the /usr/src/shadow-YYMMDD/config.h file.

If you define this constant and then compile, you must create an /etc/gshadow file to hold the group passwords and the group administrator information.

When you created the /etc/shadow, you used a program called pwconv, there no equivalent program to create the /etc/gshadow file, but it really doesn’t matter, it takes care of itself.

To create the initial /etc/gshadow file do the following:

     touch /etc/gshadow
     chown root.root /etc/gshadow
     chmod 700 /etc/gshadow

Once you create new groups, they will be added to the /etc/group and the /etc/gshadow files. If you modify a group by adding or removing users or changing the group password, the /etc/gshadow file will be changed.

The programs groups, groupadd, groupmod, and groupdel are provided as part of the Shadow Suite to modify groups.

The format of the /etc/group file is as follows:

     groupname:!:GID:member,member,...

Where:

groupname

The name of the group

!

The field that normally holds the password, but that is now relocated to the /etc/gshadow file.

GID

The numerical group ID number

member

List of group members The format of the /etc/gshadow file is as follows:

     groupname:password:admin,admin,...:member,member,...

Where:

groupname

The name of the group

password

The encoded group password.

admin

List of group administrators

member

List of group members The command gpasswd is used only for adding or removing administrators and members to or from a group. root or someone in the list of administrators may add or remove group members. The groups password can be changed using the passwd command by root or anyone listed as an administrator for the group. Despite the fact that there is not currently a manual page for gpasswd, typing gpasswd without any parameters gives a listing of options. It’s fairly easy to grasp how it all works once you understand the file formats and the concepts.

7.5 Consistency checking programs

7.5.1 pwck The program pwck is provided to provide a consistency check on the /etc/passwd and /etc/shadow files. It will check each username and verify that it has the following:

• the correct number of fields

• unique user name

• valid user and group identifier

• valid primary group

• valid home directory

• valid login shell

It will also warn of any account that has no password.

It’s a good idea to run pwck after installing the Shadow Suite. It’s also a good idea to run it periodically, perhaps weekly or monthly. If you use the -r option, you can use cron to run it on a regular basis and have the report mailed to you.

7.5.2 grpck

grpck is the consistency checking program for the /etc/group and /etc/gshadow files. It performs the following checks:

• the correct number of fields

• unique group name

• valid list of members and administrators

It also has the -r option for automated reports.

7.6 Dial-up passwords.

Dial-up passwords are another optional line of defense for systems that allow dial-in access. If you have a system that allows many people to connect locally or via a network, but you want to limit who can dial in and connect, then dial-up passwords are for you. To enable dial-up passwords, you must edit the file

/etc/login.defs and ensure that DIALUPS CHECK ENAB is set to yes.

Two files contain the dial-up information, /etc/dialups which contains the ttys (one per line, with the leading ”/dev/” removed). If a tty is listed then dial-up checks are performed.

The second file is the /etc/d passwd file. This file contains the fully qualified path name of a shell, followed by an optional password.

If a user logs into a line that is listed in /etc/dialups, and his shell is listed in the file /etc/d passwd he will be allowed access only by suppling the correct password.

Another useful purpose for using dial-up passwords might be to setup a line that only allows a certain type of connect (perhaps a PPP or UUCP connection). If a user tries to get another type of connection (i.e. a list of shells), he must know a password to use the line.

Before you can use the dial-up feature, you must create the files.

The command dpasswd is provided to assign passwords to the shells in the /etc/d passwd file. See the manual page for more information.

8 Adding shadow support to a C program

Adding shadow support to a program is actually fairly straightforward. The only problem is that the program must be run by root (or SUID root) in order for the the program to be able to access the /etc/shadow file.

This presents one big problem: very careful programming practices must be followed when creating SUID programs. For instance, if a program has a shell escape, this must not occur as root if the program is SUID root.

For adding shadow support to a program so that it can check passwords, but otherwise does need to run as root, it’s a lot safer to run the program SUID shadow instead. The xlock program is an example of this.

In the example given below, pppd-1.2.1d already runs SUID as root, so adding shadow support should not make the program any more vulnerable.

8.1 Header files

The header files should reside in /usr/include/shadow. There should also be a /usr/include/shadow.h,

but it will be a symbolic link to /usr/include/shadow/shadow.h.

To add shadow support to a program, you need to include the header files:

#include <shadow/shadow.h>
#include <shadow/pwauth.h>

It might be a good idea to use compiler directives to conditionally compile the shadow code (I do in the example below).

8.2 libshadow.a library

When you installed the Shadow Suite the libshadow.a file was created and installed in /usr/lib.

When compiling shadow support into a program, the linker needs to be told to include the libshadow.a library into the link.

This is done by:

     gcc program.c -o program -lshadow

However, as we will see in the example below, most large programs use a Makefile, and usually have a variable called LIBS=... that we will modify.

8.3 Shadow Structure

The libshadow.a library uses a structure called spwd for the information it retrieves from the /etc/shadow file. This is the definition of the spwd structure from the /usr/include/shadow/shadow.h header file:

struct spwd
{
  char *sp_namp;                /* login name */
  char *sp_pwdp;                /* encrypted password */
  sptime sp_lstchg;             /* date of last change */
  sptime sp_min;                /* minimum number of days between changes */
  sptime sp_max;                /* maximum number of days between changes */
  sptime sp_warn;               /* number of days of warning before password
                                   expires */
  sptime sp_inact;              /* number of days after password expires
                                   until the account becomes unusable. */
  sptime sp_expire;             /* days since 1/1/70 until account expires
*/
  unsigned long sp_flag;        /* reserved for future use */
};

The Shadow Suite can put things into the sp pwdp field besides just the encoded passwd. The password field could contain:

     username:Npge08pfz4wuk;@/sbin/extra:9479:0:10000::::

This means that in addition to the password, the program /sbin/extra should be called for further authentication. The program called will get passed the username and a switch that indicates why it’s being called. See the file /usr/include/shadow/pwauth.h and the source code for pwauth.c for more information.

What this means is that we should use the function pwauth to perform the actual authentication, as it will take care of the secondary authentication as well. The example below does this.

The author of the Shadow Suite indicates that since most programs in existence don’t do this, and that it may be removed or changed in future versions of the Shadow Suite.

8.4 Shadow Functions

The shadow.h file also contains the function prototypes for the functions contained in the libshadow.a library:

extern void setspent __P ((void));
extern void endspent __P ((void));
extern struct spwd *sgetspent __P ((__const char *__string));
extern struct spwd *fgetspent __P ((FILE *__fp));
extern struct spwd *getspent __P ((void));
extern struct spwd *getspnam __P ((__const char *__name));
extern int putspent __P ((__const struct spwd *__sp, FILE *__fp));

The function that we are going to use in the example is: getspnam which will retrieve for us a spwd structure for the supplied name.

8.5 Example

This is an example of adding shadow support to a program that needs it, but does not have it by default.

This example uses the Point-to-Point Protocol Server (pppd-1.2.1d), which has a mode in which it performs PAP authentication using user names and passwords from the /etc/passwd file instead of the PAP or CHAP files. You would not need to add this code to pppd-2.2.0 because it’s already there.

This feature of pppd probably isn’t used very much, but if you installed the Shadow Suite, it won’t work anymore because the passwords are no longer stored in /etc/passwd.

The code for authenticating users under pppd-1.2.1d is located in the /usr/src/pppd-1.2.1d/pppd/auth.c file.

The following code needs to be added to the top of the file where all the other #include directives are. We have surrounded the #includes with conditional directives (i.e. only include if we are compiling for shadow support).

#ifdef HAS_SHADOW
#include <shadow.h>
#include <shadow/pwauth.h>
#endif

The next thing to do is to modify the actual code. We are still making changes to the auth.c file.

Function auth.c before modifications:

/*
 * login - Check the user name and password against the system
 * password database, and login the user if OK.
 *
 * returns:
 *      UPAP_AUTHNAK: Login failed.
 *      UPAP_AUTHACK: Login succeeded.
 * In either case, msg points to an appropriate message.
 */
static int
login(user, passwd, msg, msglen)
    char *user;
    char *passwd;
    char **msg;
    int *msglen;
{
    struct passwd *pw;
    char *epasswd;
    char *tty;

if ((pw = getpwnam(user)) == NULL) {
    return (UPAP_AUTHNAK);
}
 /*
 * XXX If no passwd, let them login without one.
 */
if (pw->pw_passwd == '\0') {
    return (UPAP_AUTHACK);
}
epasswd = crypt(passwd, pw->pw_passwd);
if (strcmp(epasswd, pw->pw_passwd)) {
    return (UPAP_AUTHNAK);
}
syslog(LOG_INFO, "user %s logged in", user);
/*
 * Write a wtmp entry for this user.
 */
tty = strrchr(devname, '/');
if (tty == NULL)
    tty = devname;
else
    tty++;
logwtmp(tty, user, "");             /* Add wtmp login entry */
logged_in = TRUE;
return (UPAP_AUTHACK);
}

The user’s password is placed into pw->pw passwd, so all we really need to do is add the function getspnam. This will put the password into spwd->sp pwdp.

We will add the function pwauth to perform the actual authentication. This will automatically perform secondary authentication if the shadow file is setup for it.

Function auth.c after modifications to support shadow:

/*
* login - Check the user name and password against the system
* password database, and login the user if OK.
*
* This function has been modified to support the Linux Shadow Password
* Suite if USE_SHADOW is defined.

*
* returns:
*      UPAP_AUTHNAK: Login failed.
*      UPAP_AUTHACK: Login succeeded.
* In either case, msg points to an appropriate message.
*/
static int
login(user, passwd, msg, msglen)
char *user;
char *passwd;
char **msg;
int *msglen;
{
struct passwd *pw;
char *epasswd;
char *tty;
#ifdef USE_SHADOW
struct spwd *spwd;
struct spwd *getspnam();
#endif
if ((pw = getpwnam(user)) == NULL) {
    return (UPAP_AUTHNAK);
}
#ifdef USE_SHADOW
    spwd = getspnam(user);
    if (spwd)
            pw->pw_passwd = spwd->sp-pwdp;
#endif
 /*
 * XXX If no passwd, let NOT them login without one.
 */
if (pw->pw_passwd == '\0') {
    return (UPAP_AUTHNAK);
}
#ifdef HAS_SHADOW
if ((pw->pw_passwd && pw->pw_passwd[0] == '@'
     && pw_auth (pw->pw_passwd+1, pw->pw_name, PW_LOGIN, NULL))
    || !valid (passwd, pw)) {
    return (UPAP_AUTHNAK);
}

#else
    epasswd = crypt(passwd, pw->pw_passwd);
    if (strcmp(epasswd, pw->pw_passwd)) {
        return (UPAP_AUTHNAK);
    }
#endif
    syslog(LOG_INFO, "user %s logged in", user);
    /*
     * Write a wtmp entry for this user.
     */
    tty = strrchr(devname, '/');
    if (tty == NULL)
        tty = devname;
    else
        tty++;
    logwtmp(tty, user, "");             /* Add wtmp login entry */
    logged_in = TRUE;
    return (UPAP_AUTHACK);
}

Careful examination will reveal that we made another change as well. The original version allowed access (returned UPAP AUTHACK if there was NO password in the /etc/passwd file. This is not good, because a common use of this login feature is to use one account to allow access to the PPP process and then check the username and password supplied by PAP with the username in the /etc/passwd file and the password

in the /etc/shadow file.

So if we had set the original version up to run as the shell for a user i.e. ppp, then anyone could get a ppp connection by setting their PAP to user ppp and a password of null.

We fixed this also by returning UPAP AUTHNAK instead of UPAP AUTHACK if the password field was empty.

Interestingly enough, pppd-2.2.0 has the same problem.

Next we need to modify the Makefile so that two things occur: USE SHADOW must be defined, and libshadow.a needs to be added to the linking process.

Edit the Makefile, and add:

     LIBS = -lshadow

Then we find the line:

     COMPILE_FLAGS = -I.. -D_linux_=1 -DGIDSET_TYPE=gid_t

And change it to:

     COMPILE_FLAGS = -I.. -D_linux_=1 -DGIDSET_TYPE=gid_t -DUSE_SHADOW

Now make and install.

9 Frequently Asked Questions.

Q: I used to control which tty’s root could log into using the file /etc/securettys, but it doesn’t seem to work anymore, what’s going on?

A: The file /etc/securettys does absolutely nothing now that the Shadow Suite is installed. The tty’s that root can use are now located in the login configuration file /etc/login.defs. The entry in this file may point to another file.

Q: I installed the Shadow Suite, but now I can’t login, what did I miss?

A: You probably installed the Shadow programs, but didn’t run pwconv or you forgot to copy /etc/npasswd

to /etc/passwd and /etc/nshadow to /etc/shadow. Also, you may need to copy login.defs to /etc.

Q: In the section on xlock, it said to change the group ownership of the /etc/shadow file to shadow. I don’t have a shadow group, what do I do?

A: You can add one. Simply edit the /etc/group file, and insert a line for the shadow group. You need to ensure that the group number is not used by another group, and you need to insert it before the nogroup entry. Or you can simply suid xlock to root.

Q: Is there a mailing list for the Linux Shadow Password Suite?

A: Yes, but it’s for the development and beta testing of the next Shadow Suite for Linux. You can get added

to the list by mailing to: shadow-list-request@neptune.cin.net with a subject of: subscribe. The list

is actually for discussions of the Linux shadow-YYMMSS series of releases. You should join if you want to get involved in further development or if you install the Suite on your system and want to get information on newer releases.

Q: I installed the Shadow Suite, but when I use the userdel command, I get ”userdel: cannot open shadow group file”, what did I do wrong?

A: You compiled the Shadow Suite with the SHADOWGRP option enabled, but you don’t have an /etc/gshadow file. You need to either edit the config.h file and recompile, or create an /etc/group file. See the section on shadow groups.

Q: I installed the Shadow Suite but now I’m getting encoded passwords back in my /etc/passwd file, what’s wrong?

A: You either enabled the AUTOSHADOW option in the Shadow config.h file, or your libc was compiled with the SAHDOW COMPAT option. You need to determine which is the problem, and recompile.

10 Copyright Message.

The Linux Shadow Password HOWTO is Copyright (c) 1996 Michael H. Jackson.

Permission is granted to make and distribute verbatim copies of this document provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this document under the conditions for verbatim copies above, provided a notice clearly stating that the document is a modified version is also included in the modified document.

Permission is granted to copy and distribute translations of this document into another language, under the conditions specified above for modified versions.

Permission is granted to convert this document into another media under the conditions specified above for modified versions provided the requirement to acknowledge the source document is fulfilled by inclusion of an obvious reference to the source document in the new media. Where there is any doubt as to what defines ’obvious’ the copyright owner reserves the right to decide.

11 Miscellaneous and Acknowledgments.

The code examples for auth.c are taken from pppd-1.2.1d and ppp-2.1.0e, Copyright (c) 1993 and The Australian National University and Copyright (c) 1989 Carnegie Mellon University.

Thanks to Marek Michalkiewicz <marekm@i17linuxb.ists.pwr.wroc.pl¿ for writing and maintaining the Shadow Suite for Linux, and for his review and comments on this document.

Thanks to Ron Tidd <rtidd@tscnet.com¿ for his helpful review and testing.

Thanks to everyone who has sent me feedback to help improve this document.

Please, if you have any comments or suggestions then mail them to me.

regards

Michael H. Jackson <mhjack@tscnet.com¿